# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 13.2.1 Energy Consumption Comparison
In contrast, PoS is designed to be far more energy-efficient. Because validators in PoS are chosen based on the amount they have staked, there is no need for computationally intensive mining. As a result, PoS can significantly reduce the environmental impact of blockchain networks. For example, Ethereum’s switch from PoW to PoS with Ethereum 2.0 is projected to reduce the energy consumption of the network by over 99% [257] .
@@ 13.2.2 Network Security Comparison
The security models of PoS and PoW differ significantly. In PoW, the security of the network relies on the computational power of miners. An attack vector in PoW is the 51% attack, where an entity that controls more than 51% of the network’s hash rate can theoretically manipulate the blockchain by double-spending or preventing transactions[258].
PoS, on the other hand, secures the network by relying on economic incentives. The primary attack vector in PoS is the ”Nothing at Stake” problem[259], where validators may attempt to validate multiple chains simultaneously without any risk. However, PoS addresses this issue by introducing penalties such as slashing[260], where dishonest validators can lose their staked funds if they try to attack the network.
@@ 13.2.3 Impact on Centralization and Decentralization
Both PoW and PoS face challenges related to centralization, but in different ways. PoW tends to favor those with access to more powerful hardware and cheaper electricity[261], leading to the rise of large mining pools that can dominate the network.
In PoS, centralization can occur due to wealth concentration[262]. Since validators are chosen based on their stake, those with more cryptocurrency have a higher chance of being selected, which can lead to the rich becoming richer. However, some PoS systems implement measures to encourage decentralization, such as limiting the influence of any single validator and ensuring that smaller validators have a chance to participate[263].
@@ 13.3 Applications of PoS
Ethereum 2.0 marks one of the most prominent examples of a major blockchain transitioning from PoW to PoS[230]. In Ethereum 2.0, participants must lock up at least 32 ETH to become a validator. Validators are responsible for proposing new blocks and validating transactions, and they are rewarded for doing so correctly.
Ethereum 2.0 introduces sharding[232], which divides the blockchain into smaller parts called ”shards” to improve scalability. By combining PoS with sharding, Ethereum aims to increase the number of transactions it can handle while also significantly reducing its energy consumption compared to PoW.
@@ 13.3.2 Other Blockchain Projects Using PoS
  • Cardano: Cardano[264] uses a PoS protocol called Ouroboros, where validators (or stake pool operators) are randomly chosen to validate blocks. Cardano emphasizes a research-driven approach to PoS and provides a unique staking delegation model, allowing users to delegate their stake to stake pools without losing control of their funds.
  • Polkadot: Polkadot[265] employs a Nominated Proof of Stake (NPoS) system, where nominators back validators with their stake. This system promotes decentralization by distributing staking power among multiple validators.
  • Tezos: Tezos[266] uses a variation of PoS called Liquid Proof of Stake (LPoS), where validators, known as ”bakers,” are selected to validate transactions. Tezos also allows users to delegate their staking power without locking up their funds, providing flexibility to participants.
@@ 13.4.1 The “Nothing at Stake” Problem
One of the potential issues in PoS is the ”Nothing at Stake” problem[259]. Since validators in PoS do not need to expend significant resources to validate blocks, they might be tempted to validate multiple chains, hoping that one will eventually be accepted by the network. This could undermine the security of the blockchain.
@@ 13.4.2 Long-Range Attacks and Their Prevention
Another challenge in PoS is long-range attacks[267], where an attacker with control of old validator keys attempts to rewrite the blockchain history. These attacks are particularly concerning for PoS because validators are chosen based on stake, and old keys could be used to manipulate the chain.
To prevent long-range attacks, blockchain systems employ strategies such as checkpointing[268]. Checkpoints are specific points in the blockchain that are considered final, and the history before these checkpoints cannot be altered. This helps protect the blockchain from attackers attempting to rewrite older blocks.
@@ 14.1 Basic Working Principles of DPoS
Delegated Proof of Stake (DPoS) is a consensus mechanism that builds upon the traditional Proof of Stake (PoS) system[269], designed to achieve a higher level of scalability and efficiency while maintaining security. DPoS introduces the concept of delegating power to a smaller group of trusted participants, referred to as delegate nodes, to validate transactions and maintain the blockchain. These delegates are chosen through a voting process, which is an integral part of the DPoS framework. This section will discuss the fundamental principles that define how DPoS works, focusing on the voting mechanism and the responsibilities of elected delegate nodes.
@@ 14.1.1 Voting Mechanism for Electing Nodes
In DPoS, the stakeholders of the network, often referred to as token holders, have the ability to vote for a specific number of delegates[270]. These delegates, also known as block producers or witnesses, are responsible for validating transactions, producing blocks, and securing the network.
@@
Example: In the EOS blockchain[271, 272], there are 21 active block producers at any given time. Every 126 seconds, new block producers are elected based on stakeholder votes. This means that if a block producer does not perform well, such as by failing to produce blocks on time or acting maliciously, they can be quickly replaced through a democratic voting process.
@@ 14.1.2 Responsibilities of Delegate Nodes
    Transaction Validation: Delegate nodes are responsible for verifying and validating transactions. Once validated, these transactions are grouped into blocks, which are then added to the blockchain. Since only a limited number of delegates are elected, this process is highly efficient, leading to faster transaction confirmation times compared to traditional PoS or Proof of Work (PoW) systems[272].
@@
    Block Production: In DPoS, the elected delegates take turns producing blocks. Each delegate is assigned a specific time slot, and during this slot, they must produce a block[273]. If a delegate fails to produce a block during their slot, the network may skip that block, potentially impacting transaction throughput.
@@
    Accountability: One of the critical features of DPoS is the constant accountability of delegate nodes[274]. If a delegate node underperforms, such as by not producing blocks or engaging in malicious behavior, stakeholders can vote to remove them from their position. This feature promotes an environment of trust and responsibility, ensuring that only reliable delegates remain in power.
Example: In the Steem blockchain[275], delegates (referred to as witnesses) are responsible for producing blocks every three seconds. If a witness fails to produce a block, they may be penalized, and stakeholders can vote them out of the witness position. This dynamic system helps ensure high network uptime and reliability.
@@ 14.2 Comparison Between DPoS and PoS
Delegated Proof of Stake (DPoS) and Proof of Stake (PoS) are both consensus mechanisms designed to secure blockchain networks, but they differ in how they approach decentralization, security, and efficiency[276]. This section will explore the differences between DPoS and PoS, focusing on the issues of node centralization and the balance between security and efficiency.
@@ 14.2.1 The Issue of Node Centralization
One of the key distinctions between DPoS and PoS is the potential for centralization[277]. In a PoS system, any stakeholder can participate in the consensus process, meaning that there could be a broad distribution of validators. The more decentralized the network is, the less likely it is to suffer from a concentration of power, making the network more resistant to corruption or manipulation.
@@
Example: In EOS, only 21 block producers are responsible for maintaining the network, which can lead to a concentration of power among a few delegates[278]. While these delegates are held accountable by the voting system, the limited number of participants means that the network may be more centralized compared to a PoS system with hundreds or thousands of validators.
@@ 14.2.2 Balancing Security and Efficiency
DPoS is designed to balance security and efficiency, focusing on maximizing transaction throughput while maintaining a certain level of security. By reducing the number of nodes involved in the block validation process, DPoS can achieve much higher transaction speeds than PoS or PoW systems[279].
@@
Example: In BitShares[280], the DPoS system allows the network to process thousands of transactions per second, making it highly efficient for financial applications. However, stakeholders must remain vigilant to ensure that block producers are acting in the network’s best interest to avoid security risks.
@@ 14.3 Applications of DPoS
Several blockchain platforms use DPoS to enhance scalability and transaction throughput. This section will examine how DPoS is implemented in prominent platforms like EOS[271], Steem[275], and BitShares[280].
@@ 14.3.1 DPoS Mechanism in EOS
EOS is one of the most well-known platforms using the DPoS consensus mechanism[271]. It achieves high transaction speeds by allowing only 21 elected block producers to validate transactions. These block producers are voted in by EOS token holders and can be replaced if they do not perform their duties effectively.
@@ 14.3.2 DPoS Structure in Steem and BitShares
Steem[275] and BitShares[280] also implement DPoS but with slightly different structures and approaches. In Steem, elected witnesses are responsible for producing blocks and managing the network’s governance. Similarly, in BitShares, the DPoS mechanism ensures high transaction throughput while providing a decentralized voting system for governance decisions.
@@ 15.1 Proof of Authority (PoA)
Proof of Authority (PoA)[281] is a consensus mechanism that relies on a limited number of trusted validators, who are responsible for validating transactions and creating new blocks. This system is well-suited for private or consortium blockchains where decentralization is not the main concern, but trust, speed, and low cost are prioritized.
@@ 15.1.1 Working Principles of PoA
Proof of Authority operates differently from mechanisms like Proof of Work (PoW) and Proof of Stake (PoS). In PoA, a set of pre-approved validators with established identities is responsible for maintaining the integrity of the network[281]. Unlike PoW, where computational power is required to validate transactions, or PoS, where validators stake cryptocurrency, PoA leverages the reputation and authority of known validators. The validators are typically well-known and trusted entities, such as companies or individuals with a public reputation.
The process of validating transactions and producing blocks in PoA networks is fast because the small number of validators reduces the complexity involved[282]. The validators are incentivized to maintain the network’s trust, as any malicious behavior could harm their reputation, leading to their removal from the role of validator.
@@ 15.1.2 Validator Selection and Network Governance
In a PoA network, the process of selecting validators is critical. Validators are typically selected based on their trustworthiness and reputation[283]. In many cases, validators are well-established companies or organizations that have an interest in maintaining the stability and security of the network.
One of the key governance mechanisms in PoA systems is transparency[284]. Since the identities of the validators are known, the community can hold them accountable for their actions. Additionally, the governance system may include rules that allow for the replacement of validators if they act maliciously or lose the trust of the community.
@@ 15.1.3 Applications of PoA
  • Private Blockchains: Many enterprises use PoA for their private blockchains[285]. These blockchains are closed to the public, and only pre-approved participants (validators) are allowed to take part in the network. The fast transaction speed and low costs make PoA ideal for supply chain management and other enterprise applications.
  • Consortium Blockchains: In consortium blockchains, multiple organizations come together to form a blockchain network. Validators are selected from the member organizations, and trust is maintained through pre-existing business relationships[283]. An example of this is VeChain, a blockchain designed for supply chain management that uses a PoA consensus mechanism.
@@ 15.2 Hybrid Consensus Mechanisms
Hybrid consensus mechanisms combine different consensus algorithms to balance the trade-offs in security, energy efficiency, and decentralization[286]. The most common hybrid models combine Proof of Work (PoW) and Proof of Stake (PoS), where each system contributes its strengths to the overall network.
@@ 15.3 Emerging Consensus Mechanisms
As blockchain technology evolves, new consensus mechanisms are being developed to address limitations in scalability, security, and privacy. Two such mechanisms are Practical Byzantine Fault Tolerance (PBFT)[287] and the integration of Zero-Knowledge Proofs (ZKPs)[288] in consensus systems.
@@ 15.3.1 Practical Byzantine Fault Tolerance (PBFT)
PBFT is a consensus algorithm that is designed to handle Byzantine faults in a distributed system[289]. Byzantine faults occur when some participants in the system act maliciously or send conflicting information. PBFT achieves consensus through multiple rounds of voting between participants, ensuring that even if some participants are compromised, the network can still reach a consensus.
@@ 15.3.2 Zero-Knowledge Proofs and Consensus Mechanisms
Zero-Knowledge Proofs (ZKPs) [290]are cryptographic techniques that allow one party to prove to another party that a statement is true without revealing any information beyond the fact that the statement is true. ZKPs are increasingly being explored as a way to enhance privacy and scalability in blockchain systems.
@@ 16.1 Decentralized Finance (DeFi)
Decentralized Finance, commonly referred to as DeFi, is an emerging financial ecosystem built on blockchain technology. It seeks to replicate and enhance traditional financial services by removing intermediaries such as banks, brokers, and financial institutions[291]. Instead of relying on centralized entities, DeFi utilizes blockchain to create an open and permissionless financial network that anyone with an internet connection can access.

### v3

@@ 13.2.1 Energy Consumption Comparison
In contrast, PoS is designed to be far more energy-efficient. Because validators in PoS are chosen based on the amount they have staked, there is no need for computationally intensive mining. As a result, PoS can significantly reduce the environmental impact of blockchain networks. For example, Ethereum’s switch from PoW to PoS with Ethereum 2.0 is projected to reduce the energy consumption of the network by over 99% [Kapengut2022AnES] .
@@ 13.2.2 Network Security Comparison
The security models of PoS and PoW differ significantly. In PoW, the security of the network relies on the computational power of miners. An attack vector in PoW is the 51% attack, where an entity that controls more than 51% of the network’s hash rate can theoretically manipulate the blockchain by double-spending or preventing transactions[Sayeed2019AssessingBC].
PoS, on the other hand, secures the network by relying on economic incentives. The primary attack vector in PoS is the ”Nothing at Stake” problem[Li2017SecuringPB], where validators may attempt to validate multiple chains simultaneously without any risk. However, PoS addresses this issue by introducing penalties such as slashing[Roughgarden2024KeynotePS], where dishonest validators can lose their staked funds if they try to attack the network.
@@ 13.2.3 Impact on Centralization and Decentralization
Both PoW and PoS face challenges related to centralization, but in different ways. PoW tends to favor those with access to more powerful hardware and cheaper electricity[Romiti2019ADD], leading to the rise of large mining pools that can dominate the network.
In PoS, centralization can occur due to wealth concentration[He2020StakingPC]. Since validators are chosen based on their stake, those with more cryptocurrency have a higher chance of being selected, which can lead to the rich becoming richer. However, some PoS systems implement measures to encourage decentralization, such as limiting the influence of any single validator and ensuring that smaller validators have a chance to participate[Li2023RewardDI].
@@ 13.3 Applications of PoS
Ethereum 2.0 marks one of the most prominent examples of a major blockchain transitioning from PoW to PoS[Asif2023ShapingTF]. In Ethereum 2.0, participants must lock up at least 32 ETH to become a validator. Validators are responsible for proposing new blocks and validating transactions, and they are rewarded for doing so correctly.
Ethereum 2.0 introduces sharding[CortesGoicoechea2020ResourceAO], which divides the blockchain into smaller parts called ”shards” to improve scalability. By combining PoS with sharding, Ethereum aims to increase the number of transactions it can handle while also significantly reducing its energy consumption compared to PoW.
@@ 13.3.2 Other Blockchain Projects Using PoS
  • Cardano: Cardano[cardanoHomeCardanoorg] uses a PoS protocol called Ouroboros, where validators (or stake pool operators) are randomly chosen to validate blocks. Cardano emphasizes a research-driven approach to PoS and provides a unique staking delegation model, allowing users to delegate their stake to stake pools without losing control of their funds.
  • Polkadot: Polkadot[pokladotPolkadotInteroperable] employs a Nominated Proof of Stake (NPoS) system, where nominators back validators with their stake. This system promotes decentralization by distributing staking power among multiple validators.
  • Tezos: Tezos[tezosHomeTezos] uses a variation of PoS called Liquid Proof of Stake (LPoS), where validators, known as ”bakers,” are selected to validate transactions. Tezos also allows users to delegate their staking power without locking up their funds, providing flexibility to participants.
@@ 13.4.1 The “Nothing at Stake” Problem
One of the potential issues in PoS is the ”Nothing at Stake” problem[Li2017SecuringPB]. Since validators in PoS do not need to expend significant resources to validate blocks, they might be tempted to validate multiple chains, hoping that one will eventually be accepted by the network. This could undermine the security of the blockchain.
@@ 13.4.2 Long-Range Attacks and Their Prevention
Another challenge in PoS is long-range attacks[Deirmentzoglou2019ASO], where an attacker with control of old validator keys attempts to rewrite the blockchain history. These attacks are particularly concerning for PoS because validators are chosen based on stake, and old keys could be used to manipulate the chain.
To prevent long-range attacks, blockchain systems employ strategies such as checkpointing[Azouvi2022PikachuSP]. Checkpoints are specific points in the blockchain that are considered final, and the history before these checkpoints cannot be altered. This helps protect the blockchain from attackers attempting to rewrite older blocks.
@@ 14.1 Basic Working Principles of DPoS
Delegated Proof of Stake (DPoS) is a consensus mechanism that builds upon the traditional Proof of Stake (PoS) system[Saad2021ComparativeAO], designed to achieve a higher level of scalability and efficiency while maintaining security. DPoS introduces the concept of delegating power to a smaller group of trusted participants, referred to as delegate nodes, to validate transactions and maintain the blockchain. These delegates are chosen through a voting process, which is an integral part of the DPoS framework. This section will discuss the fundamental principles that define how DPoS works, focusing on the voting mechanism and the responsibilities of elected delegate nodes.
@@ 14.1.1 Voting Mechanism for Electing Nodes
In DPoS, the stakeholders of the network, often referred to as token holders, have the ability to vote for a specific number of delegates[Misic2024TowardDI]. These delegates, also known as block producers or witnesses, are responsible for validating transactions, producing blocks, and securing the network.
@@
Example: In the EOS blockchain[eosHomeEOSIO, Benhaim2021ScalingBC], there are 21 active block producers at any given time. Every 126 seconds, new block producers are elected based on stakeholder votes. This means that if a block producer does not perform well, such as by failing to produce blocks on time or acting maliciously, they can be quickly replaced through a democratic voting process.
@@ 14.1.2 Responsibilities of Delegate Nodes
    Transaction Validation: Delegate nodes are responsible for verifying and validating transactions. Once validated, these transactions are grouped into blocks, which are then added to the blockchain. Since only a limited number of delegates are elected, this process is highly efficient, leading to faster transaction confirmation times compared to traditional PoS or Proof of Work (PoW) systems[Benhaim2021ScalingBC].
@@
    Block Production: In DPoS, the elected delegates take turns producing blocks. Each delegate is assigned a specific time slot, and during this slot, they must produce a block[Wang2020RevisitingTF]. If a delegate fails to produce a block during their slot, the network may skip that block, potentially impacting transaction throughput.
@@
    Accountability: One of the critical features of DPoS is the constant accountability of delegate nodes[Li2023ResearchAI]. If a delegate node underperforms, such as by not producing blocks or engaging in malicious behavior, stakeholders can vote to remove them from their position. This feature promotes an environment of trust and responsibility, ensuring that only reliable delegates remain in power.
Example: In the Steem blockchain[steemPoweringCommunities], delegates (referred to as witnesses) are responsible for producing blocks every three seconds. If a witness fails to produce a block, they may be penalized, and stakeholders can vote them out of the witness position. This dynamic system helps ensure high network uptime and reliability.
@@ 14.2 Comparison Between DPoS and PoS
Delegated Proof of Stake (DPoS) and Proof of Stake (PoS) are both consensus mechanisms designed to secure blockchain networks, but they differ in how they approach decentralization, security, and efficiency[SchaafAnalysisOP]. This section will explore the differences between DPoS and PoS, focusing on the issues of node centralization and the balance between security and efficiency.
@@ 14.2.1 The Issue of Node Centralization
One of the key distinctions between DPoS and PoS is the potential for centralization[article]. In a PoS system, any stakeholder can participate in the consensus process, meaning that there could be a broad distribution of validators. The more decentralized the network is, the less likely it is to suffer from a concentration of power, making the network more resistant to corruption or manipulation.
@@
Example: In EOS, only 21 block producers are responsible for maintaining the network, which can lead to a concentration of power among a few delegates[eosnetworkConsensusDeveloper]. While these delegates are held accountable by the voting system, the limited number of participants means that the network may be more centralized compared to a PoS system with hundreds or thousands of validators.
@@ 14.2.2 Balancing Security and Efficiency
DPoS is designed to balance security and efficiency, focusing on maximizing transaction throughput while maintaining a certain level of security. By reducing the number of nodes involved in the block validation process, DPoS can achieve much higher transaction speeds than PoS or PoW systems[Bachani2022PreferentialDP].
@@
Example: In BitShares[bitsharesBitSharesBlockchain], the DPoS system allows the network to process thousands of transactions per second, making it highly efficient for financial applications. However, stakeholders must remain vigilant to ensure that block producers are acting in the network’s best interest to avoid security risks.
@@ 14.3 Applications of DPoS
Several blockchain platforms use DPoS to enhance scalability and transaction throughput. This section will examine how DPoS is implemented in prominent platforms like EOS[eosHomeEOSIO], Steem[steemPoweringCommunities], and BitShares[bitsharesBitSharesBlockchain].
@@ 14.3.1 DPoS Mechanism in EOS
EOS is one of the most well-known platforms using the DPoS consensus mechanism[eosHomeEOSIO]. It achieves high transaction speeds by allowing only 21 elected block producers to validate transactions. These block producers are voted in by EOS token holders and can be replaced if they do not perform their duties effectively.
@@ 14.3.2 DPoS Structure in Steem and BitShares
Steem[steemPoweringCommunities] and BitShares[bitsharesBitSharesBlockchain] also implement DPoS but with slightly different structures and approaches. In Steem, elected witnesses are responsible for producing blocks and managing the network’s governance. Similarly, in BitShares, the DPoS mechanism ensures high transaction throughput while providing a decentralized voting system for governance decisions.
@@ 15.1 Proof of Authority (PoA)
Proof of Authority (PoA)[ethereumProofofauthorityPoA] is a consensus mechanism that relies on a limited number of trusted validators, who are responsible for validating transactions and creating new blocks. This system is well-suited for private or consortium blockchains where decentralization is not the main concern, but trust, speed, and low cost are prioritized.
@@ 15.1.1 Working Principles of PoA
Proof of Authority operates differently from mechanisms like Proof of Work (PoW) and Proof of Stake (PoS). In PoA, a set of pre-approved validators with established identities is responsible for maintaining the integrity of the network[ethereumProofofauthorityPoA]. Unlike PoW, where computational power is required to validate transactions, or PoS, where validators stake cryptocurrency, PoA leverages the reputation and authority of known validators. The validators are typically well-known and trusted entities, such as companies or individuals with a public reputation.
The process of validating transactions and producing blocks in PoA networks is fast because the small number of validators reduces the complexity involved[Islam2022ACA]. The validators are incentivized to maintain the network’s trust, as any malicious behavior could harm their reputation, leading to their removal from the role of validator.
@@ 15.1.2 Validator Selection and Network Governance
In a PoA network, the process of selecting validators is critical. Validators are typically selected based on their trustworthiness and reputation[ZHANG2019181]. In many cases, validators are well-established companies or organizations that have an interest in maintaining the stability and security of the network.
One of the key governance mechanisms in PoA systems is transparency[fastercapitalProofAuthority]. Since the identities of the validators are known, the community can hold them accountable for their actions. Additionally, the governance system may include rules that allow for the replacement of validators if they act maliciously or lose the trust of the community.
@@ 15.1.3 Applications of PoA
  • Private Blockchains: Many enterprises use PoA for their private blockchains[Samuel2021ChoiceOE]. These blockchains are closed to the public, and only pre-approved participants (validators) are allowed to take part in the network. The fast transaction speed and low costs make PoA ideal for supply chain management and other enterprise applications.
  • Consortium Blockchains: In consortium blockchains, multiple organizations come together to form a blockchain network. Validators are selected from the member organizations, and trust is maintained through pre-existing business relationships[ZHANG2019181]. An example of this is VeChain, a blockchain designed for supply chain management that uses a PoA consensus mechanism.
@@ 15.2 Hybrid Consensus Mechanisms
Hybrid consensus mechanisms combine different consensus algorithms to balance the trade-offs in security, energy efficiency, and decentralization[KumarJha2023HybridCM]. The most common hybrid models combine Proof of Work (PoW) and Proof of Stake (PoS), where each system contributes its strengths to the overall network.
@@ 15.3 Emerging Consensus Mechanisms
As blockchain technology evolves, new consensus mechanisms are being developed to address limitations in scalability, security, and privacy. Two such mechanisms are Practical Byzantine Fault Tolerance (PBFT)[Li2020ASM] and the integration of Zero-Knowledge Proofs (ZKPs)[ZHOU2024103678] in consensus systems.
@@ 15.3.1 Practical Byzantine Fault Tolerance (PBFT)
PBFT is a consensus algorithm that is designed to handle Byzantine faults in a distributed system[Sakho2020ResearchOA]. Byzantine faults occur when some participants in the system act maliciously or send conflicting information. PBFT achieves consensus through multiple rounds of voting between participants, ensuring that even if some participants are compromised, the network can still reach a consensus.
@@ 15.3.2 Zero-Knowledge Proofs and Consensus Mechanisms
Zero-Knowledge Proofs (ZKPs) [Berentsen2023AnIT]are cryptographic techniques that allow one party to prove to another party that a statement is true without revealing any information beyond the fact that the statement is true. ZKPs are increasingly being explored as a way to enhance privacy and scalability in blockchain systems.
@@ 16.1 Decentralized Finance (DeFi)
Decentralized Finance, commonly referred to as DeFi, is an emerging financial ecosystem built on blockchain technology. It seeks to replicate and enhance traditional financial services by removing intermediaries such as banks, brokers, and financial institutions[Gupta2022DecentralizedFA]. Instead of relying on centralized entities, DeFi utilizes blockchain to create an open and permissionless financial network that anyone with an internet connection can access.